Calculate absolute potential at 25 cm distance from 5 C charge.


1
Expert's answer
2022-06-06T08:53:41-0400
"V=k\frac{q}{r}=9*10^9*\frac{5}{0.25}=1.8*10^{11}\:\rm V"